We sell our product for $20 per unit, variable costs of $10 and fixed costs of $200,000. If we sell 25,000 units, how much is our margin of safety in dollars?
Answer #1

Contribution margin per unit = Selling price per unit - Variable cost per unit

= 20 - 10 = $10

Breakeven point in units = Fixed cost/Contribution margin per unit

= 200,000/10 = 20,000 units

Breakeven point in dollars = 20,000 units * 20 = 400,000

Margin of Safety in dollars

= Sales revenue - Breakeven point in dollars

= (20*25,000) - 400,000

= 500,000 - 400,000

= $100,000